OpenStudy (anonymous):

Write an equation that represents the problem, then solve the problem. Seven-fourths decreased by a number is 2. Find the number 7/4 - x = 2 x = -1/4. 1/4 (7-4 x) = 2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

The equation is right ! It is : \[\frac74-x=2\] Now you have to solve it !

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

Or : \[\frac{-1}4\] ??

OpenStudy (anonymous):

-1/4

OpenStudy (anonymous):

i dont get ut

OpenStudy (anonymous):

The equation is : \[\frac74-x=2\] So : \[-x=2-\frac74\] so : \[-x=\frac14\] so : \[x=\frac{-1}4\]